OpenStudy (anonymous):

Help me solve this question: The sum of three consecutive integers is 72. What are the integers.

OpenStudy (chriss):

x+x+1+x+2=72 3x+3=72 3x=69 x=23 23+24+25=72

OpenStudy (anonymous):

Thank you :)

OpenStudy (chriss):

yw :)